英語 (English)

主題別英語 (Thematic English) (工 ((生 A)1 年))

マーシェソジェラルド · 准教授/ 鳴門教育大学 1 単位後期金 5 ・ 6

(平成 19 年度以前の授業科目: 『主題別英語』 ) (平成 16 年度以前 (医保は 17 年度以前) の授業科目: 『英語 (2)』 )

【授業の目的】This class will focus on primarily on listening, speaking and writing.Students will complete a variety of integrated tasks to improve their General English ability

【授業の概要】 This class will require students to complete theme-based integrated tasks. Some themes will be general, such as ‘taking vacations’, and other themes will center around the current events and will be decided later (depending on what students wish and what is happening in the news, etc.)

【キーワード】 [キーワード] 【先行科目】 [先行科目] 【関連科目】 [関連科目] 【到達目標】The purpose of this class is to improve students’ basic English

skills.The syllabus will be task based. Tasks will be integrated, targeting a variety of discrete skills. Grammar and vocabulary items will be addressed, but they will be viewed as subservient to meaning, rather than generative and devoid of context. The communicative aspect of language will be at the forefront.

【授業の計画】 1. Introductions / Getting to know your fellow students and teacher

2. Discussing language learning strategies and goals

3. Pronunciation workshop

4. How to take point form notes effectively

5. Using past tense to talk about events

6. Language for cause and effect, drawing conclusions

7. Expressing cause and effect in writing, developing a written argument

8. Preparation for presentations

9. Presentations

10. Using adjectives and descriptive language to describe cities and places

11. Describing your hometown, describing other cities

12. Pragmatically correct usage of various past tenses

13. Describing a vacation you have taken

14. Test practice 15. Test

16. Final class 【教科書】 Students will use material prepared by the instructor and self-generated

material. There will be no textbook. 【参考書等】 [参考資料] 【成績評価の方法】 Presentation: 25 Homework assignments: 25 Speaking test: 25

Participation and attendance: 25 【再試験の有無】 Students who miss the exam or fail to perform adequately will

not receive a reexamination except under special circumstances. Sufficient documentation must be provided.

【受講者へのメッセージ】 This is a general English course and will focus on many English skills in an integrated way. You have already learned a lot of grammar but most of you have not had much opportunity to use what you have learned. In this class, you will be required to actually use the English which you have learned in the past. You have all got a lot of English ’packed’ inside your head. My job as your teacher is to help you get it out; to give you opportunities to speak. I will teach grammar and vocabulary sometimes, but not usually. Many grammar and vocabulary points will come up in the lesson, as you complete work and talk with your classmates. It is very important to ask questions in class, since this will raise important grammar and vocabulary points. Good luck, work hard and have fun

【授業コンテンツ】 http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=221193 【連絡先 (オフィスアワー・研究室・ E メールアドレス)】 ⇒ マーシェソ .

English

Thematic English (工 ((生 A)1 年))

Gerard Marchesseau · Associate Professor / Naruto University of Education 1 unit 後期金 5 ・ 6

(平成 19 年度以前の授業科目: 『主題別英語』 ) (平成 16 年度以前 (医保は 17 年度以前) の授業科目: 『英語 (2)』 )

Target⟩ This class will focus on primarily on listening, speaking and writing.Students will complete a variety of integrated tasks to improve their General English ability

Outline⟩ This class will require students to complete theme-based integrated tasks. Some themes will be general, such as ‘taking vacations’, and other themes will center around the current events and will be decided later (depending on what students wish and what is happening in the news, etc.)

�������⟩ [キーワード] Fundamental Lecture⟩ [先行科目] Relational Lecture⟩ [関連科目] Goal⟩ The purpose of this class is to improve students’ basic English skills.The

syllabus will be task based. Tasks will be integrated, targeting a variety of discrete skills. Grammar and vocabulary items will be addressed, but they will be viewed as subservient to meaning, rather than generative and devoid of context. The communicative aspect of language will be at the forefront.

Schedule⟩

1. Introductions / Getting to know your fellow students and teacher

2. Discussing language learning strategies and goals

3. Pronunciation workshop

4. How to take point form notes effectively

5. Using past tense to talk about events

6. Language for cause and effect, drawing conclusions

7. Expressing cause and effect in writing, developing a written argument

8. Preparation for presentations

9. Presentations

10. Using adjectives and descriptive language to describe cities and places

11. Describing your hometown, describing other cities

12. Pragmatically correct usage of various past tenses

13. Describing a vacation you have taken

14. Test practice 15. Test

16. Final class Textbook⟩ Students will use material prepared by the instructor and self-generated

material. There will be no textbook. Reference⟩ [参考資料] Evaluation Criteria⟩ Presentation: 25 Homework assignments: 25 Speaking test:

25 Participation and attendance: 25 Re-evaluation⟩ Students who miss the exam or fail to perform adequately will

not receive a reexamination except under special circumstances. Sufficient documentation must be provided.

Message⟩ This is a general English course and will focus on many English skills in an integrated way. You have already learned a lot of grammar but most of you have not had much opportunity to use what you have learned. In this class, you will be required to actually use the English which you have learned in the past. You have all got a lot of English ’packed’ inside your head. My job as your teacher is to help you get it out; to give you opportunities to speak. I will teach grammar and vocabulary sometimes, but not usually. Many grammar and vocabulary points will come up in the lesson, as you complete work and talk with your classmates. It is very important to ask questions in class, since this will raise important grammar and vocabulary points. Good luck, work hard and have fun

Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=221193 Contact (Office-Hour, Room, E-mail)⟩

⇒ Marchesseau .